OpenStudy (anonymous):

identify the 35th term of an arithmetric sequence where a1= -7 and a18= 95 please ?

OpenStudy (anonymous):

i have no idea how to do it, im confused

OpenStudy (anonymous):

See , read the question and tell me what is given.

OpenStudy (anonymous):

a1 and a18 are given

OpenStudy (anonymous):

Yes Apply formula for them and show me

OpenStudy (anonymous):

i dont understand what n is

OpenStudy (anonymous):

The 1 and 18 are n

OpenStudy (anonymous):

Just substitute 18 in place of n in the formula I gave above and tell me what you get

OpenStudy (anonymous):

a18 =18 ( r n=1 ?

OpenStudy (anonymous):

See @CCarmichael18 , you have substituted 18 in place of a . I am asking you to substituted 18 in place of n

OpenStudy (anonymous):

a18= a(r18-1?

OpenStudy (anonymous):

Yes What is 18-1 ??

OpenStudy (anonymous):

@shivam_bhalla ,brother its an arithmetic sequence not goemetric

OpenStudy (anonymous):

17 ?

OpenStudy (anonymous):

. *

OpenStudy (anonymous):

Extremely sorry Siddharh

OpenStudy (anonymous):

@shivam_bhalla have you been showing me the wrong thing ?

OpenStudy (anonymous):

@CCarmichael18 Unfortunately yes. You have to use this formula \[\huge a_n = a+(n-1)d\] d-->common difference = a_2 - a_1 = a_3 - a_2

OpenStudy (anonymous):

Try again. Put 18 in place of n and tell me what you get

OpenStudy (anonymous):

a18 = a+ (18-1) d ?

OpenStudy (anonymous):

a18= a(17) d

OpenStudy (anonymous):

You forgot the + sign It should be a18= a + (17) d Also value of a18 is given in the question. Substitute that and tell me what you get

OpenStudy (anonymous):

95 = a+ (17) d ?

OpenStudy (anonymous):

isnt a = -7 ?

OpenStudy (anonymous):

Yes @CCarmichael18 , was just about to tell you

OpenStudy (anonymous):

Now solve and tell me d = ?

OpenStudy (anonymous):

95 = -7+ (17) d ? d= 6?

OpenStudy (anonymous):

Great. Now what we have to find in the question.??

OpenStudy (anonymous):

the 35th term

OpenStudy (anonymous):

Yes. tell me what you will do now ?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

See just substitute n=35 in the formula above . substitute your d and a too and tell me what you get

OpenStudy (anonymous):

Any problem @CCarmichael18 ??

OpenStudy (anonymous):

yes i still dont get it im doing something wrong

OpenStudy (anonymous):

Ok. Go step by step. Tell me what you get when you put n=35 in the formula above

OpenStudy (anonymous):

85= -7 + (35-1) d ?

OpenStudy (anonymous):

How did you get 85?? It should be a_35

OpenStudy (anonymous):

a18= 95 ?

OpenStudy (anonymous):

Our task is to find a35

OpenStudy (anonymous):

so 35 = -7 + ( 35-1) d?

OpenStudy (anonymous):

Not exactly ,It should be a35 = -7 + ( 35-1) d And we also found d. Now substitute d and tell me a35=?

OpenStudy (anonymous):

197 ?

OpenStudy (anonymous):

Exactly! you found out the answer . Good work :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome :)